# ./references.bib

### v1

@@ 4 Quadratic composite operator
Suppose that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian, and some $\epsilon\geq 0$ we define the non-Hermitian $\epsilon$-pseudospectrum as
$\Lambda_{\epsilon}^{\text{Q}}\left(\bm{A},\bm{B}\right)=\left\{(\bm{\lambda},\bm{\nu})\in\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}\mid\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)\leq\epsilon\right\}.$
If we have $\epsilon=0$ in the definition of the $\epsilon$-pseudospectrum, there is a joint eigenstate $\bm{\psi}$ of $A_{i}$, and $B_{j}$ for all $i$ and $j$. This is more of a mathematical statement as we know that in physical systems, observables/operators do not commute and thus do not have a joint eigenspace. From this point of view, the Proposition 4.2 is a more useful result for physical systems when $\epsilon>0$.
In the study of physical systems we would like to see how the quadratic gap function behaves as we perturb our system as well as when we perturb the probe site. When we treat the quadratic gap as a function of the probe site alone, we can think of the probe site as taking a guess at a point in the approximate joint spectrum of the matrices $A_{i}$ and $B_{i}$. When we treat the quadratic gap as a function of the input matrices alone, we can think of studying perturbations to physical systems such as introducting point defects, line defects, etc. Thus, a useful property of the quadratic gap is lipschitz continuity in $((\bm{A},\bm{B}),(\bm{\lambda},\bm{\nu}))$. For one this implies that the quadratic gap function is differentiable almost everwhere in the measure theory sense. It also means that if we are running numerical simulations, determining a grid size for the probe sites can be predetermined ahead of time. In the case where only the probe site varies, machine precision $\epsilon_{\text{mach}}$ implies that between grid spaces, the precision is also $\epsilon_{\text{mach}}$.
@@
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{Q}}\left(\bm{C},\bm{D}\right)\right|\leq\left\lVert(\bm{A},\bm{B})-(\bm{C},\bm{D})\right\rVert+\left\lVert(\bm{\lambda},\bm{\nu})-(\bm{\alpha},\bm{\beta})\right\rVert.$
@@
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{Q}}\left(\bm{E},\bm{F}\right)\right|$ $=\left|\sigma_{\min}\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)-\sigma_{\min}\left(M_{(\bm{\alpha},\bm{\beta})}\left(\bm{E},\bm{F}\right)\right)\right|.$
Combine this with perturbation of singular values on page 78 of [bhatia] and we are done. ∎
@@ 5 Relationship between Clifford linear, radial and quadratic gaps
As mentioned in Section 3 we noted that the Clifford composite operator is non-normal. This means that the variation in the spectrum can vary drastically.
Despite this we have that Theorem 1 in [ruhe],
@@
By claim (ii) in [kahan], equivalence of norms and by a geometric argument we have that
@@
Indeed this formally tells us that if $\nu$ is real valued and $B$ is Hermitian, the two gaps are the same. In fact $B$ Hermitian and $\nu$ real means we have equality with the spectral localizer gap in [cerjan_loring_vides_2023], and [loring].
@@
$\left|\dot{\mu}_{(\bm{\lambda},\nu)}^{\text{C}}\left(\bm{A},B\right)-\mu_{(\bm{\lambda},\nu)}^{\text{Q}}\left(\bm{A},B\right)\right|\leq\sqrt{\sum_{i<j}||[A_{i},A_{j}]||+||F||},$
@@
$\sqrt{\sigma_{\min}\left(\text{Q}_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)}=\sqrt{\sigma_{\min}\left(K\right)}.$
@@
Proposition 5.2 is also equivalent Proposition II.4 in [cerjan_loring_vides_2023] when the the operator $B$ is Hermitian and the probe site $\nu$ for be is restricted to be real valued.
@@ 6 Local nature of the quadratic composite operator
The idea here with the non-Hermitian composite operator being local is that if we had a physical system, a perturbation to the Hamiltonian far away from the probe site $(\bm{\lambda},\bm{\nu})$ will have little to no effect on the quadratic gap. In other words small defects away from a probe site should not have significantly large effects when probing away from the defect. The following theorem establishes this idea. In fact this is a generalization of Theorem III.1 from [cerjan_loring_vides_2023], and the proof in in the same spirit.
For the remainder of the section we will assume that the entries of the matrix $d_{1}$ tuple in $\bm{A}$ commute and that each of the $A_{i}$ is invertible. In a physical setting where $A_{i}=X_{i}$ are position observables this is easily accomplished by shifting the system in position so that the corresponding operator is invertible. Because of the linear properties of the quadratic operator we consider $(\bm{\lambda},\bm{\nu})=\vec{\bm{0}}$. This allows us to consider the following construction. It can be though of as the Euclidian distance from the origin in spacial coordinates. Specifically,
$Z=\sqrt{A_{1}^{2}+\cdots+A_{d_{1}}^{2}}$
@@
$\left\lVert Z^{-1}\left(\sum_{j=1}^{d_{2}}\left(B_{j}^{\dagger}C_{i}+C_{i}^{\dagger}B_{j}+C_{i}^{\dagger}C_{i}\right)\right)Z^{-1}\right\rVert\leq C$
for some $C<1$, then
$(1-C)^{1/2}\mu_{\vec{\bm{0}}}^{\text{Q}}\left(\bm{A},\bm{B}\right)\leq\mu_{\vec{\bm{0}}}^{\text{Q}}\left(\bm{A},\bm{B}+\bm{C}\right)\leq(1+C)^{1/2}\mu_{\vec{\bm{0}}}^{\text{Q}}\left(\bm{A},\bm{B}\right).$
@@
$-CZ^{2}\leq\sum_{j=1}^{d_{2}}\left(B_{j}^{\dagger}C_{i}+C_{i}^{\dagger}B_{j}+C_{i}^{\dagger}C_{i}\right)\leq CZ^{2}.$
@@
$\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)$ $\leq Z^{2}+Y^{2}+CZ^{2}$
$\leq(1+C)\left(Z^{2}+Y^{2}\right)$
$=(1+C)\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)$
$\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)$ $\geq Z^{2}+Y^{2}-CZ^{2}$
$\geq(1-C)\left(Z^{2}+Y^{2}\right)$
$=(1-C)\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right).$
@@
$(1+C)^{-1}\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\leq\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\leq(1-C)^{-1}\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}.$
@@
$(1+C)^{-1}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert\leq\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\right\rVert\leq(1-C)^{-1}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert.$
@@
$(1-C)^{\frac{1}{2}}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert^{-\frac{1}{2}}\leq\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\right\rVert^{-\frac{1}{2}}\leq(1+C)^{\frac{1}{2}}\left\lVert\text{Q}_{\vec{\bm{0}}}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert^{-\frac{1}{2}}.$
Note that in the proof we make extensive use that for positive semidefinite matrices $A,B$, $A\geq B$ if and only if $A-B$ is positive semidefinite. This generates an ordering on the set of positive semidefinite matrices.
@@ 7 Applications to a point gaped system
In order to encounter non-Hermiticity in physical systems one need not go very far. Indeed we consider a two level system such as the one mentioned in [tls_paper]. The two level system gives rise to a non-Hermitian Hamiltonian operator. In particular we have a point gap system. We will provide some numerical justification for why the Clifford radial gap might be the best one to use over the established Clifford linear gap. To illustrate, we first consider a general case with shifted energy as follows:
@@
with $\Delta E,\Delta\gamma,c\in\mathbb{R}$. By doing a traditional eigenenergy and eigenstate analysis we have that the eigenenergy and eigenstate pair are
@@
If we consider the case when $\Delta E=0$ then we have an exceptional point when $c=\mp\Delta\gamma/2$. In this scenario we have the eigenenergy and eigenstate pair
@@
In other words the eigenspace of $H$ becomes degenerate! The quadratic gap for this setup can get complicated, so we compute one of the exceptional point cases when $21=c=\Delta\gamma/$ 2. We will fix the position probe site of $Q$ to be $x=0$ so that the non-Hermitian Clifford and quadratic composite operators become
$L_{(0,E)}(X,H)=\begin{bmatrix}2i-E&1&-1&0\\ 1&-E&0&1\\ -1&0&-2i+\overline{E}&-1\\ 0&1&-1&-\overline{E}\\ \end{bmatrix},\mbox{ }Q_{(0,E)}(X,H)=\begin{bmatrix}\left|E\right|^{2}+4\textrm{Im}(E)+6&-2\textrm{Re}(E)-2i\\ -2\textrm{Re}(E)+2i&2+\left|E\right|^{2}\end{bmatrix}$
In Table 1 we compute the Clifford linear gap, radial gap, the quadratic gap, as well as the differences between each gap. We can see that the Clifford radial gap (Figure 2) detects the exceptional point at probe site $E=i$ and agrees well with the quadratic gap (Figure 3) near the exceptional point well. This can be seen by looking at the difference of the two gaps in Figure 6. On the other hand, while the Clifford linear gap agrees with the quadratic gap near the exceptional point (Figure 5), by itself it does not detect well the exceptional point (Figure 1). This is the motivation for defining the so called Clifford radial gap.
@@
Table 1: Shown are the Clifford linear and radial gaps (Figure 1 and 2), quadratic gap (Figure 3) as well as the difference between the gaps in Figure 4 5, 6 of a Two Level System with $c=1$,$\Delta\gamma=2$ and position probe site $x=0$.

### v2

@@ 4 Quadratic composite operator
Suppose that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian, and some $\epsilon\geq 0$ we define the non-Hermitian right, left, and quadratic $\epsilon$-pseudospectrum as
$\Lambda_{\epsilon}^{\text{RQ}}\left(\bm{A},\bm{B}\right)$ $=\left\{(\bm{\lambda},\bm{\nu})\in\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}\mid\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}\right)\leq\epsilon\right\}$
$\Lambda_{\epsilon}^{\text{LQ}}\left(\bm{A},\bm{B}\right)$ $=\left\{(\bm{\lambda},\bm{\nu})\in\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}\mid\mu_{(\bm{\lambda},\bm{\nu})}^{\text{LQ}}\left(\bm{A},\bm{B}\right)\leq\epsilon\right\}$
$\Lambda_{\epsilon}^{\text{Q}}\left(\bm{A},\bm{B}\right)$ $=\left\{(\bm{\lambda},\bm{\nu})\in\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}\mid\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)\leq\epsilon\right\}$
If we have $\epsilon=0$ in the definition of the $\epsilon$-pseudospectrum, there is a joint eigenvector $\bm{\psi}$ of $A_{i}$, and $B_{j}$ for all $i$ and $j$. This is more of a mathematical statement as we know that in physical systems, observables/operators do not commute and thus do not have a joint eigenspace. From this point of view, the Proposition 4.4 is a more useful result for physical systems when $\epsilon>0$.
In the study of physical systems we would like to see how the quadratic gap function behaves as we perturb our system as well as when we perturb the probe site. When we treat the quadratic gap as a function of the probe site alone, we can think of the probe site as taking a guess at a point in the approximate joint spectrum of the matrices $A_{i}$ and $B_{i}$. When we treat the quadratic gap as a function of the input matrices alone, we can think of studying perturbations to physical systems such as introducting point defects, line defects, etc. Thus, a useful property of the right quadratic gap is Lipschitz continuity in $((\bm{A},\bm{B}),(\bm{\lambda},\bm{\nu}))$. It also means that if we are running numerical simulations, determining a grid size for the probe sites can be predetermined ahead of time. In the case where only the probe site varies, machine precision $\epsilon_{\text{mach}}$ implies that between grid spaces, the precision is also $\epsilon_{\text{mach}}$.
@@
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{RQ}}\left(\bm{C},\bm{D}\right)\right|\leq\left\lVert(\bm{A},\bm{B})-(\bm{C},\bm{D})\right\rVert+\left\lVert(\bm{\lambda},\bm{\nu})-(\bm{\alpha},\bm{\beta})\right\rVert.$
@@
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{RQ}}\left(\bm{E},\bm{F}\right)\right|$ $=\left|\sigma_{\min}\left(RM_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)-\sigma_{\min}\left(RM_{(\bm{\alpha},\bm{\beta})}\left(\bm{E},\bm{F}\right)\right)\right|.$
Combine this with perturbation of singular values on page 78 of [27] and we are done. ∎
Corollary 4.9.
$\left|\mu_{(\bm{\lambda},\bm{\nu})}^{\text{LQ}}\left(\bm{A},\bm{B}\right)-\mu_{(\bm{\alpha},\bm{\beta})}^{\text{LQ}}\left(\bm{C},\bm{D}\right)\right|\leq\left\lVert(\bm{A},\bm{B}^{\dagger})-(\bm{C},\bm{D}^{\dagger})\right\rVert+\left\lVert(\bm{\lambda},\bm{\nu}^{*})-(\bm{\alpha},\bm{\beta}^{*})\right\rVert.$
With $\bm{B}^{\dagger}$ applying pointwise adjoint and $\bm{\nu}^{*}$ also conjugate applying pointwise to the entries.
Proof.
This follow from the fact that $LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=RQ_{(\bm{\lambda},\bm{\nu}^{*})}\left(\bm{A},\bm{B}^{\dagger}\right)$
These two results tell us that the quadratic gap function is pointwise continuous, but we conjecture that it might also be Lipschitz continuous.
Corollary 4.10.
Suppose that $\bm{A},\bm{C}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C}),\bm{B},\bm{D}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i},C_{i}$ Hermitian, $B_{j},D_{j}$ non-Hermitian, and $(\bm{\lambda},\bm{\nu}),(\bm{\alpha},\bm{\beta})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then the quadratic gap function is pointwise continuous as a function from $(\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}})\times\Pi_{i=1}^{d_{1}+d_{2}}\mathcal{M}_{n}(\mathbb{C})$ to $\mathbb{R}$.
Proof.
Lipschitz continuity of the the left and right quadratic operator gap functions as well as pointwise continuity of the minimum function. ∎
@@ 5 Relationship between Clifford linear, radial and quadratic gaps
As mentioned in Section 3 we noted that the non-Hermitian spectral localizer is non-normal. This means that the variation in the spectrum can vary drastically.
Despite this we have that Theorem 1 in [30],
@@
By claim (ii) in [37], equivalence of norms and by a geometric argument we have that
@@
Indeed this formally tells us that if $\nu$ is real valued and $B$ is Hermitian, the two gaps are the same. In fact $B$ Hermitian and $\nu$ real means we have equality with the spectral localizer gap in [4], and [32].
@@
[…]
@@
$\sqrt{\sigma_{\min}\left(Q_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)}=\sqrt{\sigma_{\min}\left(K\right)}.$
@@
Proposition 5.2 is equivalent Proposition II.4 in Ref [4] when the the operator $B$ is Hermitian and the probe site $\nu$ is restricted to be real valued.
@@ 6 Local nature of the quadratic composite operator
The non-Hermitian composite operator being local in this setting means that if we had a physical system, a perturbation to the Hamiltonian far away from the probe site $(\bm{\lambda},\bm{\nu})$ will have little to no effect on the quadratic gap. In other words, small defects away from a probe site should not have significantly large effects when probing away from the defect. The following theorem establishes this idea. In fact this is a generalization of Theorem III.1 from [4], and the proof is in the same spirit.
For the remainder of the section we will assume that the entries of the matrix $d_{1}$ tuple in $\bm{A}$ commute and that each of the $A_{i}$ is invertible. In a physical setting where $A_{i}=X_{i}$ are position observables this is easily accomplished by shifting the system in position so that the corresponding operator is invertible.
This allows us to consider the following construction. It can be thought of as the Euclidean distance from the origin in spacial coordinates. Specifically,
$Z^{2}=\sum_{i=1}^{d_{1}}(A_{i}-\lambda_{i})^{2}$
@@
$\left\lVert Z^{-1}\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})^{\dagger}C_{j}+C_{j}^{\dagger}(B_{j}-\nu_{j})+C_{j}^{\dagger}C_{j}\right)Z^{-1}\right\rVert\leq K$
for some $K<1$, then
$(1-K)^{1/2}\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}\right)\leq\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}+\bm{C}\right)\leq(1+K)^{1/2}\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}\right).$
@@
$KZ^{2}\leq\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})^{\dagger}C_{j}+C_{j}^{\dagger}(B_{j}-\nu_{j})+C_{j}^{\dagger}C_{j}\right)\leq KZ^{2}.$
@@
$\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}+\bm{C}\right)$ $\leq Z^{2}+Y^{2}+KZ^{2}$
$=(1+K)\left(Z^{2}+Y^{2}\right)$
$=(1+K)\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$
$\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}+\bm{C}\right)$ $\geq Z^{2}+Y^{2}-KZ^{2}$
$\geq(1-K)\left(Z^{2}+Y^{2}\right)$
$=(1-K)\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right).$
@@
$(1+K)^{-1}\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)^{-1}\leq\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\leq(1-K)^{-1}\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)^{-1}.$
@@
$(1+K)^{-1}\left\lVert\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert\leq\left\lVert\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\right\rVert\leq(1-K)^{-1}\left\lVert\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert.$
@@
$(1-K)^{\frac{1}{2}}\left\lVert\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert^{-\frac{1}{2}}\leq\left\lVert\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}+\bm{C}\right)^{-1}\right\rVert^{-\frac{1}{2}}\leq(1+K)^{\frac{1}{2}}\left\lVert\text{RQ}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)^{-1}\right\rVert^{-\frac{1}{2}}.$
Note that in the proof we make extensive use that for Hermitian matrices $A,B$, $A\geq B$ if and only if ABA-B is positive semidefinite. This generates an ordering on the set of positive semidefinite matrices.
Remark 6.2.
We can also show that the left quadratic gap function is local by the same argument.
Corollary 6.3.
$\left\lVert Z^{-1}\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})^{\dagger}C_{j}+C_{j}^{\dagger}(B_{j}-\nu_{j})+C_{j}^{\dagger}C_{j}\right)Z^{-1}\right\rVert\leq K$
for some $K<1$, then
$(1-K)^{1/2}\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)\leq\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}+\bm{C}\right)\leq(1+K)^{1/2}\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right).$
Proof.
Combine Theorem 6.2, Remark 6.2 and the fact that the minimum function preserves inequalities. ∎
@@ 7 Applications to a point gaped system
In order to encounter non-Hermiticity in physical systems one need not go very far. Indeed we consider a two level system such as the one mentioned in [28]. The two level system gives rise to a non-Hermitian Hamiltonian operator. In particular we have a point gap system. We will provide some numerical justification for why the Clifford radial gap might be the best one to use over the established Clifford linear gap. To illustrate, we first consider a general case with shifted energy as follows:
@@
with $\Delta E,\Delta\gamma,c\in\mathbb{R}$. By doing a traditional eigenenergy and eigenvector analysis we have that the eigenenergy and eigenvector pair are
@@
If we consider the case when $\Delta E=0$ then we have an exceptional point when $c=\mp\Delta\gamma/2$. In this scenario we have the eigenenergy and eigenvector pair
@@
In other words the eigenspace of $H$ becomes degenerate! The quadratic gap for this setup can get complicated, so we compute one of the exceptional point cases when $1=c=\Delta\gamma/2$. We will fix the position probe site of $Q$ to be $x=0$ so that the non-Hermitian Clifford and quadratic composite operators become
$L_{(0,E)}(X,H)$ $=\begin{bmatrix}2i-E&1&-1&0\\ 1&-E&0&1\\ -1&0&-2i+E^{*}&-1\\ 0&1&-1&+E^{*}\\ \end{bmatrix}$
$Q_{(0,E)}(X,H)$ $=\begin{bmatrix}\left|E\right|^{2}+6+4i\textrm{Im}(E)&-2\textrm{Re}(E)-2i&&\\ -2\textrm{Re}(E)+2i&2+\left|E\right|^{2}&&\\ &&\left|E\right|^{2}+6+4i\textrm{Im}(E)&-2\textrm{Re}(E)+2i\\ &&-2\textrm{Re}(E)-2i&2+\left|E\right|^{2}\end{bmatrix}$
In Table 1 we compute the Clifford linear gap, radial gap, the quadratic gap, as well as the differences between each gap. We can see that the Clifford radial gap (Figure 2) detects the exceptional point at probe site $E=i$ and agrees well with the quadratic gap (Figure 3) near the exceptional point well. This can be seen by looking at the difference of the two gaps in Figure 6. On the other hand, while the Clifford linear gap agrees with the quadratic gap near the exceptional point (Figure 5), by itself it does not detect well the exceptional point (Figure 1). This is the motivation for defining the Clifford radial gap.
@@
[…]